Menu

University Libraries

All systems are online.

Maintenance for Databases Unresolved

[5/23 @ 8:26 am - Sara]

SRDS will be down for maintenance starting friday May 26 and may be down through Monday May 29

This Maintenance is scheduled from 08:26am 5/23/17 until 07:00pm 12/31/69.

Update for Catalog Resolved

[5/17 @ 3:13 pm - Ben]

We added 7,476 new Naxos MARC records today.

This Update happened on 5/17/17.

Outage for Databases Resolved

[5/17 @ 1:53 pm - Sara]

SRDS is experiencing unexpected downtime. Vendor alerted us to the situation today. They are working to restore access. Updates when available.

[5/18 @ 11:42 am - Sara]

They now expect SRDS to remain down throughout the 18th. We will update when we hear more.

[5/19 @ 1:42 pm - Sara]

Will be down the rest of 05/19. Possibly longer

[5/22 @ 7:45 am - Sara]

issue was finally resolved over the weekend. But Saturday was still reporting downtime. Possibly resolved on the 21st

[7:45 am - Sara]

This Outage was reported on 5/17/17 and resolved on 5/22/17.

Minor Issue for Databases Resolved

[5/10 @ 8:38 am - Ben]

It seems there is an issue with proxy and Safari and Android browsers requiring a certificate to get to the proxy log in page. Mary is looking into this now.

[10:00 am - Ben]

It appears IT has corrected this issue.

This Minor Issue was reported on 5/10/17 and resolved on 5/10/17.

Update for Catalog Resolved

[5/9 @ 3:22 pm - Sara]

May Serial Solutions MARC Records 05/09/2017: We added 709 new serial solutions journal Marc records today. We now have 69,009 total in our collection. ​

This Update happened on 5/9/17.

Update for Catalog Resolved

[5/5 @ 11:59 am - Ben]

We added 219 new Safari MARC records and deleted 72.

This Update happened on 5/5/17.

Update for Catalog Resolved

[5/5 @ 11:58 am - Ben]

We added 69 new NBER MARC records today.

This Update happened on 5/5/17.

Update for Catalog Resolved

[5/5 @ 9:47 am - Ben]

We added 2,167 new Kanopy MARC records today.

This Update happened on 5/5/17.

Update for Catalog Resolved

[5/5 @ 9:02 am - Ben]

We added 573 new Films on Demand MARC records today.

This Update happened on 5/5/17.

Update for Catalog Resolved

[5/5 @ 8:07 am - Ben]

We deleted 227 EBL MARC records today.

This Update happened on 5/5/17.